OpenStudy (anonymous):

-11x-10y=17 5x-7y=50 answer?

OpenStudy (anonymous):

-7*(-11x-10y=17) 10(5x-7y)=50 77x+70y=-84 50x-70y=500 +------------- 127x=416 x=416/127 50*416/127-500=70y -42700/127=70y y=-42700/8890

OpenStudy (anonymous):

x=3 y=-5

OpenStudy (anonymous):

-7*(-11x-10y=17) 10(5x-7y)=50 77x+70y=-119 50x-70y=500 +------------- 127x=381 x=381/127 x=3 50*3-500=70y -350/70=y y=-5